Find the number of Calories in a stalk of celery and in a carrot.
1. A carrot has 13 Calories more than a celery stalk.
2. Five carrots and ten celery stalks have only 170 Calories.


Answer by scott8148(6628)

let x=celery, so x+13=carrot ... 5(x+13)+10(x)=170 ... 5x+65+10x=170 ... 15x+65=170